1. What is the future value of 1000 earnings 3% per year for 50 months?
2. What is yhe present value of 1000000 receuved 50 years from now, if you wish to earn 3% per yeat interest?
3. What is the present value of 10000 per month received over the next 2 years. if you wish to each 2% interest?
4.If I deposit 5000 per year in my savings earning 2.5% interest how much would i have in 50 years?
Answer #1

Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!

Ans to 1 A=P*(1+i)^n
A= Future Value
P= 1000
i= 3%= 0.03
n= no of years= 50 months= 50/12 years
A= 1000*(1+.03)^(50/12)
A= 1000*(1.03)^(50/12)
Future Value                  1,131.07
Ans to 2 A=P*(1+i)^n
A= Future Value 1,000,000.00
P= Present Value
i= 3%= 0.03
n= no of years= 50
1000000= P*(1+.03)^50
P= 1000000
(1+.03)^50
P= 1000000
                         4.38
Present Value              228,107.08
Ans to 3 P*(1-(1/(1+r)^n)
r
P Monthly Payment        10,000.00
i Interest rate 2%
r rate per period 2/12%      0.001667
n number of periods 2 years 24 months
Present Value 10000*(1-(1/(1+.001667)^24)
0.001667
Present Value 10000*(1-(1/(1.001667)^24)
0.001667
Present Value 10000*(1-0.960813756709294)
0.001667
Present Value 10000*0.039186243290706
0.001667
Present Value 391.8624329
0.001667
Present Value              235,117.46
Ans to 4 A=P*((1+i)^n-1)
r
A= Future Value
P=                   5,000.00
i= 2.5%= 0.025
n= no of years= 50
A= 5000*((1+.025)^50-1)
0.025
A= 5000*(3.43710871970354-1)
0.025
A= 5000*2.43710871970354
0.025
A= 12185.5436
0.025
Future Value              487,421.74